OpenStudy (anonymous):

Is -7 + 9= -9 + 7 true, false, or open?

OpenStudy (anonymous):

False, -7+9=2, -9+7=-2

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

Can you help me with a couple more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Okay. Is 4x-3=19 true, false, or open?

OpenStudy (anonymous):

But not for long because I gtg in a bit

OpenStudy (anonymous):

K

OpenStudy (anonymous):

What is x?

OpenStudy (anonymous):

It doesn't say

OpenStudy (anonymous):

Oh ok, then its open

OpenStudy (anonymous):

Okay. thanks

OpenStudy (anonymous):

Anymore?

OpenStudy (anonymous):

yeah. 3 more

OpenStudy (anonymous):

not true false and open stuff tho

OpenStudy (anonymous):

try and post all of then or copy and paste then

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

It won't let me copy and paste it

OpenStudy (anonymous):

Ok type them all.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Which of the following is a solution to the question 16= 4x - 4? A. -5 B. -4 C. 5 D. 16 I'm guessing B?

OpenStudy (anonymous):

its c, 4x5=20, 20-4=16

OpenStudy (anonymous):

An art student wants to make a model of the classroom. The Length of the classroom is 2.4 times its width. The length of the student's model is 42 in. What should the width of the model be?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Hum

OpenStudy (anonymous):

Sorry i forgot the choices

OpenStudy (anonymous):

17.5 is the answer

OpenStudy (anonymous):

A. 17.5 in B. 20.5 in C. 83.6 in D. 100.8 in

OpenStudy (anonymous):

A

OpenStudy (anonymous):

It's d

OpenStudy (anonymous):

How come?

OpenStudy (anonymous):

2.4x 42=100.8

OpenStudy (anonymous):

Remember, its 2.4 should be divided, not multiplied. its 17.5

OpenStudy (anonymous):

42 should be divided by 2.4*

OpenStudy (anonymous):

im off, gtg

OpenStudy (anonymous):

k